OpenStudy (anonymous):

Earn a medal! If sin Θ = 5/6, what are the values of cos Θ and tan Θ?

OpenStudy (vishweshshrimali5):

What is the relationship between \(sin^{2}(\theta)\) and \(cos^{2}(\theta)\) ?

OpenStudy (anonymous):

The next image is the question.  First step, you solve "x" in the image. \[x^{2} + 5^{2} = 6^{2}\] \[x^{2} + 25 =36\]\[x^{2} = 36- 25 = 11\] \[x = \pm \sqrt{11}\] So the answer is \[\cos \theta = \frac{ \pm \sqrt{11} }{ 6 }\] \[\tan \theta = \frac{ 5 }{ \pm \sqrt{11} } = \pm \frac{ 5 }{ \sqrt{11} }\] If you draw the trigonometric function in the question,  Θ>90deg. is a condition of -√11.